OpenStudy (anonymous):

my brain is fried lol is decimal 2 as a fraction 0.02% or 2%

jimthompson5910 (jim_thompson5910):

you mean 0.2 right?

jimthompson5910 (jim_thompson5910):

and you want this as a fraction?

jimthompson5910 (jim_thompson5910):

if so, then 0.2 = 2/10 because the 2 is in the tenths place now you just reduce 2/10 to get your final answer

OpenStudy (anonymous):

no its 2. written as a percent sorry

jimthompson5910 (jim_thompson5910):

oh so you want 2% written in decimal form?

OpenStudy (anonymous):

yes pls

OpenStudy (anonymous):

sorry jim

jimthompson5910 (jim_thompson5910):

that's ok

jimthompson5910 (jim_thompson5910):

2% = 2/100 by definition (x% = x/100) so 2% = 2/100 = 1/50 after you reduce this means 2% converts to the fraction 1/50 ------------------------------------------ 2% = 0.02 (move the decimal 2 spots to the left) or you can just evaluate 2/100 to get 0.02 which is the same answer

jimthompson5910 (jim_thompson5910):

so 2% as a decimal is 0.02

OpenStudy (anonymous):

ok thank you, could not make my mind up if it was 0.02 or 2

jimthompson5910 (jim_thompson5910):

it's 0.02 the number 2 refers to 200%, which is quite different

OpenStudy (anonymous):

ok thank you like i said my brain fried tonite lol

jimthompson5910 (jim_thompson5910):

lol that's ok, yw

OpenStudy (anonymous):

have a good night

jimthompson5910 (jim_thompson5910):

you too